I am 18 years old. I attend the local vocational-technical school. My area of study is commercial

heating and cooling. My school tuition is relatively low and I will complete my program of study

in nine months. I can pay most of my expenses by working full-time in the summer and part-

time during the school year. I am still living at home, and I plan to get an apartment of my own

next year. I am an avid sports fan. I have a little 12-inch television. I'd like to use my credit card

to buy a flat screen 65-inch television set at a cost of $1 ,300.


1
Expert's answer
2021-04-20T18:37:33-0400

Calculate the amount of the monthly payment on the loan:

"A=S\times(r+\frac{r}{(1+r)^n-1})"

S=1300

n=9

r=0.1

"A=1300\times(0.1+\frac{0.1}{(1+0.1)^9-1})=225.73"

it is necessary to find a job with a monthly income much more2 25.73


"total refund amount=225.73\times9=2031.59"